The width of a rectangle is 5 feet. The length of the rectangle is 11 feet. What is the perimeter of the rectangle?
Answer by ReadingBoosters(3246)

p = 2(l+w) = 2l+2w
w=5
l=11
p = 2(5+11)
p=2(16)
p=32 ft